Consider the revenues and costs in 2019 for Spruce Decor​ Inc., an​ Alberta-based furniture company entirely owned by Mr. Harold Buford.
Furniture Sales
​$655 comma 000
Catalogue Sales
​$17 comma 500
Labour Costs
​$335 comma 000
Materials Costs
​$148 comma 000
Advertising Costs
​$28 comma 000
​Debt-Service Costs
​$32 comma 000

a. What would accountants determine Spruce​ Decor's profits to be in​ 2019?
Spruce​ Decor's profits in 2019 equal ​$
nothing.
b. Suppose Mr. Buford has ​$375 comma 000 of capital invested in Spruce Decor. Also suppose that equally risky enterprises earn a 16 percent rate of return on capital. What is the opportunity cost for Mr.​ Buford's ​capital?
The opportunity cost for Mr.​ Buford's capital is ​$???
nothing.
c. What are the economic profits for Spruce Decor in​ 2019?
The economic profits for Spruce Decor in 2019 equal ​$???
nothing.
1
Expert's answer
2019-11-04T09:27:19-0500

Date

(a)  What would accountants determine Spruce Decor’s profits P to be in 2010?

"P= R-C"

R= 657,000

C= 542,000

P= 115,000

(b)  Suppose Mr. Buford has $400 000 of capital invested in Spruce Decor. Also suppose that equally risky enterprises earn a 16-percent rate of return on capital. What is the opportunity cost for Mr. Buford’s capital?

"OC= 400,000*0.16"

OC=64,000

(c) What are the economic profits EP for Spruce Decor in 2010?

"EP= 115,000-64,000"

EP= 51,000